OpenStudy (anonymous):

sketch graph : f(x)=2^-x +3

OpenStudy (anonymous):

use table of functions. assume value of x to gain value of y. .

OpenStudy (anonymous):

could someone help me with my problem>

OpenStudy (anonymous):

I don't know what to do with the -x

OpenStudy (anonymous):

okay so say I said x = 1 , my y value would be 1?

OpenStudy (anonymous):

does this graph make an s shape? in a way?

OpenStudy (anonymous):

wouldn't it be 2+3? since 2^0 = 2^1 = 2?